nrestion S: A firm can use three diffrent produetion technologies, i Units of Capitat Number of Employees requirements at each level of output as follows: TechnologyUnits of Capital Number of Em 16 12 12 24 Which technology is the most labor- intensive technology? a) b) Suppose the firm is operating in a high-wage country, where capital cost is $70 per per day and labor cost is $50 per worker per day, which technology is ch e the firm is operating in a low-wage country, where capitatl per day but labor cost is only $25 per unit per day, which technology is e Now suppose unit e) rating in a low.wage country, where capital cost is s70 per hexpest d) Which technology is the most capital-intensive technology?
Answer #1

(a)

We compute the Labor/Capital (L/K = Employees/Capital) ratio for each technology as follows.

Tech A = 8/16 = 0.5

Tech B = 12/12 = 1

Tech C = 20/8 = 2.5

Tech D = 24/6 = 4

Since (L/K) ratio is highest for Technology D, this is most labor-intensive.

(b)

Total cost (C) ($) = wL + rK = 50L + 70K, computed for each technology as follows.

Tech A = 50 x 8 + 70 x 16 = 400 + 1120 = 1520

Tech B = 50 x 12 + 70 x 12 = 600 + 840 = 1440

Tech C = 50 x 20 + 8 x 70 = 1000 + 560 = 1560

Tech D = 50 x 24 + 70 x 6 = 1200 + 420 = 1620

So technology B is cheapest, its total cost being minimum.

(c)

Total cost (C) ($) = wL + rK = 25L + 70K, computed for each technology as follows.

Tech A = 25 x 8 + 70 x 16 = 200 + 1120 = 1320

Tech B = 25 x 12 + 70 x 12 = 300 + 840 = 1140

Tech C = 25 x 20 + 8 x 70 = 500 + 560 = 1060

Tech D = 25 x 24 + 70 x 6 = 600 + 420 = 1020

So technology D is cheapest, its total cost being minimum.

(d)

Capital intensity is measured as (K/L = Capital/Employees). Since (L/K) ratio is minimum (= 0.5) for technology A, (K/L) ratio is maximum for technology A, so this is the most capital intensive technology.